# Theoretical Benefit and Limitation of Diffusion Language Model

### v1

@@ 5 Experiments
Figure 1: Sampling Efficiency and Quality of MDMs on Formal Languages: The above subfigure illustrates generative perplexity of generated sequences versus the number of sampling steps for $n$-gram languages ($n\in{2,3,4}$) and HMM. The y-axis represents the generative perplexity, and the x-axis represents the sampling steps, with the last point indicating the performance of auto-regressive models. The figure below shows the SER of generated sequences versus the number of sampling steps for the same formal languages. The y-axis represents the SER, while the x-axis is the same as the above figure. The number above each bar indicates the speedup of MDMs under different sampling steps compared to the auto-regressive models.
We conducted a series of experiments to empirically validate the theoretical findings, focusing on evaluating the sampling quality and computational efficiency of MDMs under diverse metrics. The results reveal that while MDMs effectively generate low-TER sequences, but achieving low-SER demands substantial computational resources. We will first introduce our experimental settings and then present the experimental results.
@@ 5.1 Experimental Setup
Tasks and Datasets. First, we evaluated MDMs on a variety of formal languages, including $n$-gram languages ($n\in\{2,3,4\}$) and HMMs. For each formal language, parameters such as transition matrices, observation matrices, and initial distributions were generated through random sampling. Detailed descriptions of the parameter generation process, along with illustrative examples of the resulting sequences, are provided in Section E.1. Using these formal languages, we constructed datasets comprising $1{,}000{,}000$ samples, of which $990{,}000$ were allocated for training and $10{,}000$ for validation. When using the formal language models to generate the dataset, we set the max length of 512.
Figure 2: Evaluation on Language Tasks: The left subfigure illustrates the text generation quality of MDLM-OWT across different sampling steps, with GPT2-medium as baseline. The y-axis represents the average generative perplexity of 2000 generated texts, and the x-axis indicates the number of sampling steps. The numbers above indicate the speedup of MDLM-OWT under different sampling steps compared to GPT2-medium. The right subfigure shows the accuracy of MDM on the GSM8K benchmark at different sampling steps, with Qwen-Math-1.5B as baseline. The y-axis indicates accuracy, and the x-axis represents the number of sampling steps.
@@
Model Training. We adopted transformer-based architectures as the backbone models due to their scalability and expressiveness in sequence modeling tasks. Comprehensive architectural details, including the number of layers, hidden dimensions, and positional encoding schemes, are provided in Table 2 in Section E.2. The training process followed the framework proposed by Sahoo et al. (2024), with additional training configurations detailed in Table 3. Models were trained for 20 epochs, and their convergence was monitored using the validation set. Perplexity was used as the primary convergence metric, and the trained models achieved optimal perplexity values consistent with the ground-truth language models that generated the datasets.
@@
Evaluation Metrics. To assess the quality of generated sequences, we used TER and SER as the primary evaluation metrics, in alignment with our theoretical framework. Computational efficiency was evaluated based on the number of sampling steps. Following prior work (Lou et al., 2024; Xu et al., 2024), generative perplexity was employed as the TER metric to evaluate the sample qualities under different sampling steps. We compute the generative perplexity using the ground-truth model to evaluate the likelihood of sequences generated by MDMs, which were subsequently converted into perplexity scores. SER was computed directly using its definition in Equation 5, leveraging ground-truth models for evaluation. For sequence generation, we utilized the ddpm_cache sampler proposed in prior work (Sahoo et al., 2024), ensuring efficient sampling. Computational efficiency was measured by the number of sampling steps, and we further discuss the influence of ddpm_cache under different sampling steps in Appendix D. Furthermore, we also test the true speedup of MDMs under different sampling steps compared to the auto-regressive models in Figure 1, the detailed testing settings are listed in Section E.2. To ensure robust evaluation, we generated $2000$ sequences for each setting and computed both TER and SER over these samples.
@@
To compare MDMs with auto-regressive models, we trained auto-regressive models with identical architectures and model sizes on the same datasets generated by the formal languages. These models were evaluated under the same metrics, serving as a baseline for performance comparison. The training configurations are provided in Table 4.
@@ 5.2 Experiment Results
The experiment results are presented in Figure 1. The upper subfigure shows the generative perplexity across different formal languages with the number of sampling steps varying. The x-axis represents the number of sampling steps, ranging from 8 to 2048, while the y-axis measures the generative perplexity, where lower values indicate higher text fluency and token-level accuracy. The performance of auto-regressive models is marked as the final point on the x-axis for comparison. As shown in the figure, MDMs achieve near-optimal generative perplexity with relatively few sampling steps. To achieve a perplexity similar to the auto-regressive model, MDMs only require about 64 steps and demonstrate 1.57 times speedup compared to auto-regressive models. This demonstrates that MDMs are highly efficient at generating fluent sequences even with a small number of sampling steps. As the number of sampling steps increases, the performance of MDMs approaches that of auto-regressive models, converging to a similar level of generative perplexity.
The lower subfigure evaluates the relationship between the number of sampling steps and the SER, which measures the correctness of an entire sequence. The x-axis again represents the number of sampling steps, with the performance of auto-regressive models included as a baseline, and the y-axis measures the SER, where lower values indicate higher sequence-level accuracy. Compared to the upper subfigure, this subfigure reveals a slower improvement in SER as the number of sampling steps increases. For these formal languages, achieving low SER requires significantly more sampling steps. Moreover, even when the number of sampling steps reaches 2048, there remains a gap in SER between MDMs and auto-regressive models. These results demonstrate that auto-regressive models maintain a clear advantage in SER, as their token-by-token generation achieves zero SER across these tasks.
@@
Figure 1 highlights the trade-off between efficiency and accuracy for MDMs empirically. While MDMs excel in generating fluent outputs with low TER, they require substantially more sampling steps to achieve low SER, particularly for reasoning-intensive tasks that demand sequence-level correctness. These experimental results further reinforce our theoretical findings.
6 Preliminary Experiments on Large Models
We further conducted an extensive set of experiments on language tasks using open-source MDMs. First, we evaluated the quality of text generation by measuring the generative perplexity of MDMs using MDLM-OWT (Sahoo et al., 2024), a diffusion language model trained on OpenWebText (Gokaslan & Cohen, 2019). For a fair comparison, we evaluated GPT2-medium (Radford et al., 2019), which is similar in size. Second, we explored the mathematical reasoning ability of MDMs on the GSM8K dataset (Cobbe et al., 2021). Given that small models typically exhibit poor reasoning performance, we used a fine-tuned diffusion language model with 1.1B non-embedding parameters proposed by Nie et al. (2024), and compared it against model with a similar number of parameters. While the generative perplexity represents the metric of TER, mathematical reasoning is more concerned with the correctness of the entire generated sequence, thus, the GSM8K accuracy is partially consistent with the negative sequence error rate $-\operatorname{SER}$.
Text Generation. For text generation, we use MDLM-OWT, which has a context length of 1024 and size similar to GPT2-medium and is trained on OWT dataset (Gokaslan & Cohen, 2019). Since our goal is to compare the acceleration of MDMs relative to auto-regressive models and examine the effect of the number of steps on text generation quality, the absolute size and capability of the model are less important. Following the approach in the original work, we used the ddpm_cache sampler and the GPT2 tokenizer. For the number of sampling steps ranging from 4 to 2048, we generated 2000 samples of length 1024 and evaluated the generative perplexity using GPT2-large. To compare MDMs with auto-regressive models, we took GPT2-medium as baseline and computed its generative perplexity in the same manner.
The experiment result is shown in the left subfigure of Figure 2, which illustrates the text generation quality of MDLM-OWT across different sampling steps, with GPT2-medium as the baseline. The the x-axis represents the number of sampling steps, and the y-axis represents the average generative perplexity of 2000 generated texts, where lower generative perplexity indicates higher fluency and, consequently, a lower TER. The numbers above indicate the speedup of MDLM-OWT under different sampling steps compared to GPT2-medium. As is shown in the figure, MDLM-OWT matches the generative perplexity of GPT2-medium with only 32 steps, where there is a 2.28x speedup, and the perplexity continues to decline and converge as the number of sampling steps increases. This demonstrates that MDMs can generate texts efficiently while ensuring a high fluency, which illustrates the potential of MDMs for basic language generation tasks at a larger scale.
Mathematical Reasoning. For mathematical reasoning, we used the MDM provided by Nie et al. (2024), which was fine-tuned on GSM8K using a model trained on SlimPajama (Soboleva et al., 2023) for $3.3\times 10^{21}$ training FLOPs with 1.1B non-embedding parameters. This is so far the first MDM to be fine-tuned on mathematical reasoning tasks. We generated answers with a maximum length of 256 for the number of sampling steps ranging from 1 to 256. Since there are very few models fine-tuned on GSM8K at the same scale, we took Qwen2-Math-1.5B (Yang et al., 2024) as our baseline. We evaluated its performance following the widely used Language Model Evaluation Harness framework (Gao et al., 2024), and counted the average length of the generated answers, which partly reflects the efficiency of the model.
The experiment results are presented in the right subfigure of Figure 2. For all tested step numbers, the average length of generated answers for MDM is around 30, while for the baseline model, the average length is about 105. Unlike text generation, MDM does not show a significant advantage over auto-regressive models for mathematical reasoning tasks. The accuracy of the MDM decreases sharply as the number of steps falls below the sequence length, and it shows only a slight improvement as the number of samples exceeds the sequence length. While the latter may be due to the limitations of the MDM we used, the former likely caused by insufficient sampling, which leads to a high sequence error rate. It is worth noting that the experimental setup for the MDM differs from that of the baseline models, so the baseline accuracy is provided only for reference.
Summary. Figure 2 illustrates the performance of MDMs on language tasks and and its dependence on sampling steps. For text generation, MDLM-OWT achieved similar performance to GPT2-medium with few sampling steps, demonstrating efficiency in generating fluent text. On the contrary, MDMs showed no significant advantage on the GSM8K accuracy, with performance declining rapidly when the number of steps fell below the sequence length. These results highlight MDMs’ ability in text generation, but suggest challenges in reasoning-relevant tasks.
@@ 7 Conclusion and Limitations
Conclusion. This paper provides a rigorous theoretical and empirical analysis of the efficiency of MDMs under various metrics. We demonstrate that MDMs can achieve near-optimal TER with a fixed number of sampling steps, regardless of sequence length, making them highly efficient for tasks emphasizing fluency. However, when evaluated using SER, MDMs require sampling steps that scale linearly with sequence length, negating their efficiency advantage over auto-regressive models. These findings highlight the trade-off between efficiency and accuracy, depending on the evaluation metric. Experimental results further reinforce our theoretical results across formal and natural language tasks, offering practical guidance for deploying MDMs. While MDMs demonstrate efficiency advantages in applications prioritizing fluency, they may fall short in reasoning-intensive tasks requiring high accuracy compared to auto-regressive models.
Limitations. Our study focuses on formal languages modeled using HMM, which, while foundational, still differs from modern language models. Extending this analysis to more advanced language models remains an important direction for future work. Additionally, we primarily analyze Masked Diffusion Models, but the broader family of diffusion-based language models, including variants like SEDD-unform (Lou et al., 2024), requires further investigation. In summary, while our work establishes a theoretical understanding of MDMs, further exploration is needed to generalize our findings to real-world settings and to systematically analyze other diffusion approaches.
Impact Statement
This paper presents work whose goal is to advance the field of generative models. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
Appendix A Auxiliary Lemma

### v2

@@ 5 Experiments
SER with Different Sampling Steps
N-gram Length 128 256 512 1024 2048 AR
2-gram 512 $36.2\pm 4.2$ $24.5\pm 3.2$ $14.4\pm 2.9$ $7.4\pm 2.2$ $5.4\pm 1.2$ $0.2\pm 0.1$
1024 $59.5\pm 4.9$ $40.3\pm 4.6$ $23.2\pm 3.2$ $14.6\pm 3.0$ $9.4\pm 1.5$ $0.3\pm 0.2$
2048 $88.70\pm 2.8$ $71.5\pm 3.7$ $50.2\pm 4.4$ $40.2\pm 3.9$ $26.9\pm 3.3$ $0.3\pm 0.1$
3-gram 512 $37.3\pm 3.6$ $21.6\pm 3.3$ $16.4\pm 2.7$ $11.6\pm 3.0$ $7.9\pm 2.1$ $0.2\pm 0.1$
1024 $64.8\pm 4.7$ $39.4\pm 5.3$ $26.6\pm 3.0$ $19.3\pm 3.2$ $13.9\pm 2.7$ $0.3\pm 0.3$
2048 $86.0\pm 3.9$ $66.1\pm 5.3$ $48.2\pm 5.2$ $48.2\pm 5.6$ $37.8\pm 5.2$ $0.2\pm 0.2$
4-gram 512 $51.8\pm 4.4$ $39.2\pm 3.2$ $34.0\pm 3.5$ $30.0\pm 2.7$ $25.7\pm 2.6$ $0.5\pm 0.1$
1024 $67.9\pm 5.6$ $52.9\pm 5.1$ $41.8\pm 2.7$ $36.1\pm 5.8$ $35.4\pm 5.1$ $0.4\pm 0.2$
2048 $80.0\pm 4.1$ $68.1\pm 5.5$ $62.0\pm 5.4$ $62.0\pm 4.3$ $60.1\pm 3.8$ $0.5\pm 0.4$
HMM 512 $19.1\pm 2.1$ $16.7\pm 2.3$ $12.8\pm 2.6$ $9.1\pm 2.6$ $8.2\pm 2.2$ $0.4\pm 0.1$
1024 $40.1\pm 2.5$ $32.8\pm 3.1$ $29.9\pm 2.6$ $30.5\pm 3.0$ $28.0\pm 3.3$ $0.4\pm 0.2$
2048 $68.5\pm 5.6$ $63.4\pm 4.2$ $62.9\pm 5.5$ $62.9\pm 4.7$ $61.8\pm 5.8$ $0.3\pm 0.1$
TER with Different Sampling Steps
N-gram Length 128 256 512 1024 2048 AR
2-gram 512 $3.71\pm.03$ $3.68\pm.02$ $3.67\pm.01$ $3.67\pm.01$ $3.67\pm.01$ $3.67\pm.01$
1024 $3.72\pm.05$ $3.70\pm.03$ $3.68\pm.01$ $3.68\pm.02$ $3.67\pm.01$ $3.67\pm.01$
2048 $3.71\pm.05$ $3.69\pm.03$ $3.67\pm.02$ 3.66±.01plus-or-minus3.66.013.$6$01 3.66±.01plus-or-minus3.66.013.$6$01 3.66±.02plus-or-minus3.66.023.$6$02
3-gram 512 $3.13\pm.05$ $3.10\pm.03$ $3.07\pm.01$ $3.07\pm.01$ $3.07\pm.01$ $3.07\pm.01$
1024 $3.14\pm.06$ $3.10\pm.03$ $3.09\pm.02$ $3.08\pm.01$ $3.08\pm.01$ $3.08\pm.02$
2048 $3.15\pm.06$ 3.11±.03plus-or-minus3.11.033.$1$03 $3.10\pm.02$ $3.09\pm.01$ $3.08\pm.02$ $3.08\pm.02$
4-gram 512 $3.27\pm.08$ $3.23\pm.04$ $3.20\pm.02$ $3.20\pm.01$ $3.19\pm.01$ $3.19\pm.01$
1024 $3.21\pm.08$ $3.16\pm.03$ $3.16\pm.02$ $3.14\pm.01$ $3.14\pm.01$ $3.14\pm.01$
2048 $3.20\pm.06$ $3.16\pm.03$ $3.14\pm.02$ $3.13\pm.02$ $3.12\pm.01$ $3.12\pm.02$
HMM 512 $4.04\pm.03$  $4.01\pm.03$ $4.00\pm.01$ 3.99±.01plus-or-minus3.99.013.$9$01 3.99±.01plus-or-minus3.99.013.$9$01 3.99±.01plus-or-minus3.99.013.$9$01
1024 $4.03\pm.04$ $4.01\pm.02$ $4.00\pm.01$ 3.99±.02plus-or-minus3.99.023.$9$02 $3.98\pm.02$ $3.98\pm.02$
2048 $4.03\pm.02$ $4.02\pm.01$ $4.01\pm.01$ $4.00\pm.01$ $4.00\pm.01$ $4.00\pm.01$
Table 1: SER and TER results for MDMs (n-grams, $n\in\{2,3,4\}$; HMMs) vs. AR baseline across varying sequence lengths (512, 1024, 2048) and sampling steps. The ’AR’ column (underlined) provides the baseline for comparison. In the TER table, blue highlighting indicates MDM performance statistically similar to the AR baseline according to T-Test. MDMs approach AR TER with $\approx 512$ sampling steps, irrespective of sequence length. Accuracy, however, requires more steps for longer sequences and generally lags behind AR models with a large margin.
As discussed in the previous section, our theory aligns with observations in natural language tasks. However, for comprehensive validation and completeness, we also conducted experiments to empirically support our theoretical results. More precisely, these experiments were designed to investigate the relationship between sampling steps, sequence length, the SER, and the TER. We will first introduce our experimental setup, followed by a presentation of the results.
@@ 5.1 Experimental Setup
Tasks and Datasets. We evaluated MDMs on several formal languages: $n$-gram languages (with $n\in\{2,3,4\}$) and HMMs. For each language type, parameters (e.g., transition matrices, observation matrices, initial distributions) were randomly sampled. A detailed description of this generation process and examples of resulting sequences are available in Section F.1. These formal languages were used to generate datasets of $1{,}000{,}000$ samples each, with $990{,}000$ for training and $10{,}000$ for validation. Datasets were generated with sequence lengths $L\in\{512,1024,2048\}$.
@@
Model Training. Transformer-based architectures served as the backbone for our MDMs, chosen for their scalability and expressiveness in sequence modeling. Detailed architectural specifications, including layer counts, hidden dimensions, and positional encoding schemes, are provided in Table 6 (Section F.2). The training procedure largely followed the framework of Sahoo et al. (2024), with specific training configurations detailed in Table 7. Models were trained for 20 epochs, with convergence monitored on the validation set using perplexity. The trained models successfully achieved perplexity values consistent with the ground-truth language models generating the datasets.
@@
Evaluation Metrics. To evaluate the quality of generated sequences in line with our theoretical framework, we used TER and SER as primary metrics. Generative perplexity, computed using the ground-truth model, served as the TER metric. SER was calculated directly via Equation 5 (where accuracy reported in tables is $1-\text{SER}$), also utilizing the ground-truth models. For sequence generation, we employed the ddpm_cache sampler from Sahoo et al. (2024); its influence under varying sampling steps is further discussed in Section B.3. We report the inference time per sequence of MDMs with 512 sampling steps and AR models in Table 2 across different sequence length (detailed inference settings in Section F.2). For robust evaluation, all reported TER and SER values are averages over $2000$ generated sequences for each experimental setting.
@@
Auto-Regressive Baseline. We also trained AR models with identical architectures and model sizes on the same datasets generated by the formal languages for comparison. These AR models were evaluated using the same metrics. Training configurations for the AR models are detailed in Table 8.
@@ 5.2 Experiment Results
The experiment results are presented in Table 1. This table presents the generative perplexity and accuracy on n-gram languages ($n\in\{2,3,4\}$) and HMM. The results are detailed for sequence lengths of 512, 1024, and 2048, across a varying number of sampling steps (128, 256, 512, 1024, and 2048). The final ’AR’ column shows the performance of the auto-regressive baseline.
As Table 1 shows, under the metric of TER, MDMs achieve near-optimal generative perplexity, closely matching the AR baseline, with approximately 512 sampling steps. Crucially, this required number of sampling steps remains relatively constant across the tested sequence lengths (512, 1024, 2048). This empirically supports our theoretical finding that TER is primarily dependent on the number of sampling steps, rather than sequence length.
In contrast, achieving low SER with MDMs necessitates a substantially larger number of sampling steps. This requirement becomes more pronounced as sequence length increases. For instance, for 2-gram models generating sequences of length 2048, even 2048 sampling steps yield only 73.1% accuracy, whereas AR models achieve near 100%. This aligns with our theory that SER is more sensitive to both the number of sampling steps and the sequence length.
Furthermore, even with a high number of sampling steps (e.g., 2048), a significant gap in SER persists between MDMs and AR models. This is particularly evident for more complex languages like 4-grams and HMMs across all sequence lengths. For example, with 4-gram language of length 512, MDMs with 2048 steps achieve 74.3% accuracy, while the AR baseline reaches 99.5%. AR models, due to their token-by-token generation mechanism, consistently achieve near-perfect SER on these tasks. This result further validate our theoretical result that MDMs Cannot generate low-SER sentences with a low cost.
Sequence Length 512 1024 2048
MDMs (512 steps) 3.1s 4.2s 4.7s
AR 1.7s 3.3s 7.0s
Table 2: The inference time per sequence of MDMs with 512 sampling steps and AR models.
Despite the challenges in achieving low SER, MDMs demonstrate considerable efficiency in generating fluent sequences with a fixed number of sampling steps (e.g., 512 steps), especially for longer sequences. The inference time per sequence, detailed in Table 2, quantifies this advantage, showing MDMs to be highly efficient for generating fluent long sequences.
@@
These empirical findings underscores the trade-off between generation efficiency and performace for MDMs. While MDMs excel at producing fluent outputs efficiently, they require considerably more sampling iterations to achieve low SER. This characteristic is particularly salient for reasoning-intensive tasks that demand high sequence-level correctness. Collectively, these experimental results provide strong empirical corroboration for our theoretical analyses.
@@ 6 Conclusion and Limitations
This paper provides a rigorous theoretical analysis of the efficiency of MDMs under various metrics. We demonstrate that MDMs can achieve near-optimal TER with a fixed number of sampling steps, regardless of sequence length, making them highly efficient for tasks emphasizing fluency. However, when evaluated using SER, MDMs require sampling steps that scale linearly with sequence length, negating their efficiency advantage over auto-regressive models. Our study focuses on formal languages modeled using HMM, which, while foundational, still differs from modern language models. Extending this analysis to more advanced language models remains an important direction for future work. Additionally, we primarily analyze Masked Diffusion Models, but a broader family of diffusion-based language models, including variants like SEDD-unform (Lou et al., 2024). Further exploration is needed to generalize our findings to real-world settings and to systematically analyze other diffusion approaches.
@@
Appendix A Details of Preliminary Experiments
A.1 Models and Sampling Strategies Compared
We follow Ye et al. (2025) and Nie et al. (2025) to setup our experiments on GSM8K and MBPP benchmarks. For each benchmark, we compare the following sampling strategies of different models as suggested:
  • Entropy-based sampling (Ye et al., 2025) applied to Dream-v0-Base-7B, with different sampling steps.
  • Low-confidence-based sampling (Nie et al., 2025) applied to LLaDA-8B-Base, also with different sampling steps.
  • Left-to-right (AR-style) sampling strategy for Dream-v0-Base-7B and LLaDA-8B-Base, where one token is generated per step in a sequential, left-to-right manner.
  • Standard auto-regressive generation for Qwen2.5-7B with default settings.
Here, the entropy-based sampling strategy and the low-confidence-based sampling strategy are the recommended sampling strategies for the MDMs (Ye et al., 2025; Nie et al., 2025). Qwen2.5-7B serves as the standard AR baseline, while the left-to-right sampling strategy of MDMs serves as a naive baseline for the MDMs, aiming to eliminate the influence of other contributing factors, including different training datasets or training methods.
For Dream-v0-Base-7B, we manually implemented the left-to-right sampling process. In the case of LLaDA-8B-Base, setting block_length=1 during generation naturally obtains the same behavior. This configuration does not stop generation after encountering EOS, which may cause an increased time cost, but our results indicate that this behavior does not affect the overall conclusions.
A.2 Settings for GSM8k Experiments
Following Ye et al. (2025) and Nie et al. (2025), we use the widely-adopted lm-evaluation-harness [^1] https://github.com/EleutherAI/lm-evaluation-harness framework to evaluate the performance. For the GSM8K benchmark, we follow prior work and set the number of few-shot examples to 8 and the maximum length of answer to 256. We assess the GSM8K Accuracy and model efficiency of different models, where efficiency is defined by the inverse of the execution time measured on 8 Nvidia RTX 4090 GPUs with Huggingface’s transformers library [^2] https://github.com/huggingface/transformers.
In the figure, we report the relative accuracy and efficiency of MDMs compared to the baselines. These relative values are computed by dividing the actual accuracy and efficiency by those of the respective baselines – Qwen and the MDMs using the left-to-right strategy.
We list the detailed testing configurations for GSM8K in Table 3.
Models Dream w/ entropy Dream w/ left-to-right LLaDA w/ low-confidence LLaDA w/ left-to-right Qwen2.5
Max new tokens 256 256 256 256 256
Block length N/A N/A 256 1 N/A
Sampling algorithm Entropy Sequential Low-confidence Sequential N/A
Sampling steps {64,128,256} N/A {64,128,256} 256 N/A
Temperature 0 0 0 0 0
Table 3: Testing Configurations for GSM8K
A.3 Settings and Results for MBPP Experiments
For the MBPP experiment, the evaluation settings are similar to the GSM8K experiment. Specifically, use 3 few-shot examples and set the maximum answer length to 512, as suggested, to evaluate the MBPP Accuracy and efficiency. The definition of efficiency and the method for computing relative accuracy and efficiency values – by normalizing against the respective baselines – are identical to those used in the GSM8K evaluation.
The results are presented in Figure 2, and the detailed testing configurations for MBPP are listed in Table 4
Models Dream w/ entropy Dream w/ left-to-right LLaDA w/ low-confidence LLaDA w/ left-to-right Qwen2.5
Max new tokens 512 512 512 512 512
Block length N/A N/A 512 1 N/A
Sampling algorithm Entropy Sequential Low-confidence Sequential N/A
Sampling steps {128,256,512} N/A {128,256,512} 512 N/A
Temperature 0.2 0.2 0 0 0
Top p 0.95 0.95 N/A N/A N/A
Table 4: Testing Configurations for MBPP
Figure 2: Efficiency and accuracy of MDMs on MBPP (3-shot): Similar to the GSM8K experiment, the left figure uses Qwen2.5-7B as the evaluation baseline and compares the MBPP relative accuracy and efficiency of Dream-v0-Base-7B and LLaDA-8B-Base across different sampling steps. The origin can be considered as the baseline result. The first quadrant demonstrates models that can simultaneously achieve higher efficiency and accuracy. From the figure, we can see that the MDMs all fall into the third quadrant, indicating lower efficiency and lower performance. The right figure uses a baseline in which the MDMs are constrained to generate the left-most masked token at each step, effectively mimicking auto-regressive decoding. Still, even under this setting, none of the configurations fall into the first quadrant.
Appendix B Further Discussion
B.1 Various Metrics in NLP Tasks.
Evaluation metrics in NLP tasks are inherently tied to the specific objectives and requirements of their respective domains. For general language modeling tasks, perplexity (Jelinek et al., 1977; Devlin et al., 2019) remains the metric of choice due to its ability to capture a model’s predictive performance effectively. However, domain-specific tasks often demand more specialized evaluation criteria. For instance, in machine translation (Bahdanau, 2014; Wu et al., 2016), the BLEU score is widely regarded as a standard measure of translation quality (Papineni et al., 2002), while text generation tasks (Sutskever, 2014) frequently rely on metrics such as ROUGE to assess output fidelity (Lin, 2004). Similarly, tasks requiring reasoning (Wei et al., 2022c), such as mathematics (Bubeck et al., 2023) or code generation (Roziere et al., 2023; Ouyang et al., 2023), commonly adopt accuracy as an intuitive and straightforward measure of success.
B.2 Discussion Regarding the Empirical Observations.
The above results reveal that MDMs can efficiently generate low-TER sentences but may incur higher costs when evaluating the generation under SER. One might think these results are contradictory. Note that several previous works have already shown that TER (a.k.a perplexity) may not reflect a model’s true performance in solving several long-sequence understanding tasks (Huang et al., 2022; Hu et al., 2024; Luden et al., 2024). Thus, it is natural to arrive at different conclusions depending on the metric used. Moreover, many practical scenarios have shown that the choice of evaluation metric significantly influences the conclusion of other problems (see Section B.1 for various metrics in NLP tasks). For instance, while the community has previously focused on the emergence phenomenon, recent works by Wei et al. (2022a) and Schaeffer et al. (2024) demonstrate that this phenomenon may stem from the use of non-smooth evaluation metrics.
B.3 Extend to Efficient Sampling Strategies
  1. 1.
  2. 2.
Appendix C Auxiliary Lemma